RAHUL HAVE 1% SUGAR SOLUTION. HE LEFT IT ON THE ROOF AND IT BECAME 2% SOLUTION. WHAT PERCENT OF SOLUTION HAS EVAPORATED?
Answer by richwmiller(17219)

lets say he had 100 gallons
.01*100=.02(100-x)
1=2-x/50
0=1-x/50
x/50=1
x=50 gallons
50/100=1/2
1/2 evaporated
.01*y=.02(y-x)
y/100=2y/100-2x/100
y=2y-2x
2x=y
x=1/2*y
amt evaporated x is 1/2 of solution y